OpenStudy (anonymous):

Solve (2x+5)(4x−3)=0 A. x = 5/2 or x = -3/4 B. x = -5/2 or x = 3/4 C. x = 2/5 or x = 4/3 D. x = -2/5 or x = -4/3

OpenStudy (anonymous):

Are those fractions or is it division?

OpenStudy (anonymous):

fract

OpenStudy (anonymous):

2x+5=0 x=-5/2 4x-3=0 x=3/4 So, its B :)

OpenStudy (cggurumanjunath):

@LaBlueEyes2013 is it clear ?

OpenStudy (anonymous):

Well Xoya beat me to it ;) Oh well. Good job Xoya :D We've been tagging eachother in every problem I think lol. Everyone I reply to your about to reply. :P

OpenStudy (anonymous):

Hahaha I don't want to beat you :) You can answer the next one if you want I'll wait for my turn xD

OpenStudy (anonymous):

Well, don't see anything under my level. I'm in Algebra 1 so.. Lol.